# On the Formalization of Cryptographic Migration

### v3

@@ III A Model for the Cryptographic Migration Problem
All in all, post-quantum migration is the scenario that motivated our formal migration model. At the same time, it implicates several special properties which make modeling even more complex as we demonstrate in the following section.
III-A Real-World Migration Graphs
Today’s large scale IT-infrastructures are both complex and highly diverse, and examining models to create a suitable graph from a real-world IT-infrastructure is an open research topic. The aggregation of the necessary data that is used to derive the graph is highly individual and specific to the given migration goal. We state a few concrete examples of such modeling attempts:
Network migration:
When talking about network migration it seems intuitive to think of the components to represent computers, servers, and similar devices. Much thought has to be taken into the dependencies. These could, e. g. model communication paths for specific (cryptographic) protocols. For the same network, all connections between machines may be interesting regardless of the specific protocol, illustrating all the relations between machines in the network. Depending on the objective one may study several graphs representing different kinds of communication relations each – for example individual graphs for protocols like TLS/HTTPS, SSH, or IPsec/IKEv2.
System migration:
Different graphs may be modeling a specific system, e. g. the migration components being a set of software-artifacts within a firmware or operating system. It is unclear at which granularity such modeling may be done. We expect that relations may be handled similarly to SBOMs, showing dependencies between software, libraries, and so on. This may even go into the details of a specific software project showing relations between different modules, library calls, etc.
Full-scale migration:
For a system administrator, a tool focusing on the high-level network view may be enough, while programmers may profit from call-graph-like models. A full view on all abstraction levels may be provided by a combination of different graphs, e. g. a component being linked to another graph depicting the same component in a different context. This leads us to having differing graph models of the same infrastructure, just depending on our perspective.
Note that when we consider classical, naturally grown enterprise networks, the graphs will probably be both: large and complex. Also, they will have many interdependent systems. It is beyond our current model to capture time dependent changes in the migration graph. But even when we look at a snap-shot of a migration graph, the design choices for modeling real-world IT-infrastructure graphs are vast and we discovered a variety of possible ways to model such IT-infrastructures.
A reasonable assumption might be that the best strategy is to model only what is vital for the given migration and leave out all other information. Examining the exact choices, however, deserves its own research track.
III-B Real-World Patterns
$v$
$v_{1}$$v_{2}$$v_{3}$$w$
$v$$w_{2}$$w_{1}$$w_{3}$
$v$$w$
$v_{1}$$v_{2}$$v_{3}$
Figure 1: Dependency sub-graphs of real-world migration attempts: 1(a) Single component with no dependencies. 1(b) Many components depend on a single component. 1(c) One component depends on many components. 1(d) Mutual dependencies between two components. 1(e) Higher-order cycles.
When analyzing the occurrence of graph patterns in the real world, we might begin with the kinds of typical dependency patterns we observe in practice. To do so, we depicted different base-patterns in Fig. 1, from which any directed graph can be constructed. In the following, we will express real-world observations related to these patterns and discuss their semantics and meaning for our introduced theory. Also, we briefly reflect on their impact on post-quantum migration.
API-compatible change of algorithms:
When migrating an isolated (generic) technological component $v$, a simple change in the code or even an update of the relevant software migrates the component successfully. This corresponds to case Fig. 1(a). In the context of post-quantum migration this could be a simple shift from one cryptographic algorithm to an API-compatible post-quantum algorithm, e. g. the change from ECDSA to CRYSTALS-Dilithium.
Operating System dependencies:
Certain components $v_{1}$, $v_{2}$, $v_{3}$, $\dots$, for example software or libraries, can only be migrated when the underlying execution platform $w$ supports the newly migrated component. This often occurs when one version of an application is updated but the new version only runs on a newer version of the operating system. This corresponds to case Fig. 1(b). In the context of post-quantum migration, this could be the replacement of a necessary cryptographic hardware resource running a certain operating system with a PQC-ready version of it.
Regulatory requirements:
When regulatory requirements $w_{1}$, $w_{2}$, $w_{3}$, $\dots$ impose restrictions on the migratability of certain technical components $v$, only a change in the regulatory requirements enables us to actually migrate. Changing regulatory requirements, however, often depends on many further, often non-technical, components, such as laws, standardization, technological standards in general, etc. This corresponds to case Fig. 1(c) and directly applies to post-quantum migration as well.
Communication protocols:
When migrating communication protocols from one version to another, it is vital that all communicating parties in the migration cluster are migrated at the same time. This is in particular a problem when there are many parties involved and only one part of the own IT-infrastructure may be updated conveniently, e. g. by software configuration management tools with external dependencies. This problem often leads to backwards compatibility as a feature, but likely also leads to the possibility of downgrading attacks. This corresponds to case Fig. 1(d). In the context of post-quantum migration this could be the replacement of TLS by a post-quantum-ready version of it.
Dependency on space/throughput requirements:
Occasionally, there are requirements on the amount of memory, the capabilities of computing platforms, or network throughput restrictions. This corresponds to cases Figs. 1(d) and 1(e). In the context of post-quantum migration this applies when for example the key sizes of post-quantum algorithms are much larger than those of classical algorithms.
Dependencies of software-packages within distributions:
When considering software-packages and their dependencies, the graph formed this way is a directed acyclic graph with no mutual dependencies. This graph represents combinations of cases illustrated in Figs. 1(a)1(b), and 1(c). The migration of such a case can be done corresponding to the canonical topological order of this graph. In the context of post-quantum migration this could be the replacement of cryptographic libraries with post-quantum-ready versions.
Cyclically depending assets:
The draft of ETSI’s “Repeatable Framework for Quantum-safe Migrations” [6] explains that if for example three cryptographic assets are in a dependency relationship with each other, functionality issues may occur when a single asset is migrated. The functionality issues might cascade and affect all assets in the circle, even the migrated one. This corresponds to case Fig. 1(e) but the problem is mitigated in our model as we assume that such circular clusters of assets can only be migrated when all assets are migrated at the same time. In the context of post-quantum migration this might happen for example when we observe circularly cross-signed certificates. Such circular cases are rare in practice, because they create dependencies that are avoided in software architecture and engineering. Also, developers can address those kinds of dependencies by introducing adapter nodes in one direction which allow to break the dependency cycle.
We encourage the inclined reader to find more naturally occurring real-world examples. Using these base-patterns, a wide range of possible security scenarios can not only be modeled, but also described and simplified, by breaking them down into smaller, well-understood structures. Thus, once identified, the above patterns might serve as fundamental steps when deriving migration strategies in practical settings.
IV Assessing the Difficulty of the Migration Problem
We now show that, expectedly, any sufficiently large migration project has an intrinsic complexity. Specifically we show:
Theorem 1.
Expectedly we have for migration projects with $n$ migration components within our model:
  1. 1.
    The number of migration steps $M$ is expectedly $\mathsf{E}(M)\sim n/\mathop{\text{log}}\nolimits n$ with standard deviation $\mathsf{SD}(M)\sim\sqrt{n}/\mathop{\text{log}}\nolimits n$ and the expected size of each migration cluster is $\mathop{\text{log}}\nolimits n$.
  2. 2.
  3. 3.
    If the set of migratable components is at each step bound by a smooth function ${f\colon\mathbb{R}\rightarrow\mathbb{R}}$ in a real variable with $f\in o(n)$, then we expect $n/f(n)$ non-independent migration steps that successively depend on each other.
Proof.
Since every migration project decomposes disjointly into its migration clusters, we now study the expected properties of these migration clusters. Let $n$ be the number of components. There are $B_{n}$ different possible partitions of $n$ components in disjoint subsets, where $B_{n}$ is the $n$th Bell number. Denote by $M$ the number of disjoint subsets of a randomly chosen partition.
  1. 1.
    The statement on the expectation and the standard deviation of $M$ is a summary of [17] and Chapter 4, Theorem 1.1 in [22]. Furthermore, the expected maximum size of a migration cluster is, for growing $n$, asymptotic to $er-\mathop{\text{log}}\nolimits\sqrt{r}$, where $r\cdot\mathop{\text{exp}}\nolimits(r)=n$ and $r=\mathop{\text{log}}\nolimits n-(1+o(1))\mathop{\text{log}}\nolimits\mathop{\text{log}}\nolimits n\sim\mathop{\text{log}}\nolimits n$ [21]. Results on the speed of convergence are beyond the scope of this article.
  2. 2.
    Let $M_{1},\ldots,M_{s}$ be the resulting numbers of migration clusters and define $X=\mathop{\text{max}}\nolimits_{1\leq i\leq s}M_{i}-\mu$. Then conveniently $\mathsf{E}(X)=0$.
    Since for a fixed real constant $t$ the function $x\mapsto\mathop{\text{exp}}\nolimits(tx)$ is convex, by Jensen’s inequality we have:
    $\leq\sum_{i=1}^{s}\mathsf{E}(\mathop{\text{exp}}\nolimits(tM_{i}))$
    $=s\mathop{\text{exp}}\nolimits(t^{2}\sigma^{2}/2)$
    Here, the last equality follows from the fact that the moment generating function $\mathsf{E}(\mathop{\text{exp}}\nolimits(tM_{i}))$ for a normal distribution with mean zero is equal to $\mathop{\text{exp}}\nolimits(t^{2}\sigma^{2}/2)$ and in particular independent of the index-variable $i$. Rewriting this, $\mathsf{E}(X)\leq\frac{\mathop{\text{log}}\nolimits s}{t}+\frac{t\sigma^{2}}{2}$. This bound is minimal at $\frac{\sqrt{2\mathop{\text{log}}\nolimits s}}{\sigma}$ since its derivative in direction $t$ is $\frac{\sigma^{2}}{2}-\frac{\mathop{\text{log}}\nolimits s}{t^{2}}$. Thus, $\mathsf{E}(X)\leq\sigma\sqrt{2\mathop{\text{log}}\nolimits s}$ and the result follows.
  3. 3.
    We study the dependencies between the migration clusters. According to the first statement of this Theorem, we consider the directed acyclic graph resulting from the migration graph on $n$ components modulo its strongly connected components with expectedly $n^{\prime}=n/\mathop{\text{log}}\nolimits n$ vertices. Since we are considering a single migration project, we can assume that this graph is connected. We now use the standard Erdős-Rényi model [5], where this graph is constructed by connecting the components independently at random with a fixed probability $p$. One can ensure that the result is acyclic by defining any linear order on the vertices and include a directed edge only if it obeys the selected ordering. Each such edge is included in the graph with probability $p_{n}$, independently from every other edge. For such graphs, we have by Theorem 4.1 in [10], a variant of Theorem 1 in [4] that for such graphs, for growing $n$, the probability that $G$ is connected is given by
    $\begin{cases}0&\text{, if $c_{n}\to-\infty$}\\ e^{-e^{-c}}&\text{, if $c_{n}\to c$ constant}\\ 1&\text{, if $c_{n}\to\infty$}\\ \end{cases}$ (1)
    where $p_{n}=\frac{\mathop{\text{log}}\nolimits n+c_{n}}{n}$. Note that we can directly apply this Theorem to our directed acyclic graph situation, since such a graph is connected if and only if the undirected variant of it is connected.
    As our goal is to prove a lower bound on the number of edges in the condensation graph (with $n^{\prime}$ vertices!) in our model. We assume now the sparsest case $c_{n^{\prime}}\to c$ constant for which the graph is connected with high probability. Note that the doubly-exponential term in Equation (1) approaches $1$ already for comparatively small $c$ very quickly. For example, already for $c=10$, the value of $e^{-e^{-c}}\geq 0.99995$. But also for any $c>0$, we obtain for the number of edges in the condensation graph expectedly asymptotically
    $\frac{\mathop{\text{log}}\nolimits n^{\prime}+c_{n^{\prime}}}{n^{\prime}}\cdot\frac{n^{\prime}\cdot(n^{\prime}-1)}{2}\sim\frac{1}{2}n^{\prime}\mathop{\text{log}}\nolimits n^{\prime}.$
    That means that even in the sparsest case, we expect a considerable number of dependencies between the migration clusters. Furthermore, if we bound the number of components with a common ancestor of distance $k$, i. e. #$\#(v\to^{k}V)$ by a small function $f(n)=o(n)$, e. g. $f(n)=\mathop{\text{log}}\nolimits n$, we obtain by Theorem 4 in Appendix -D for the condensation graph expectedly $d(G)\geq\frac{n}{f(n)\mathop{\text{log}}\nolimits n}$, that is, large chains of dependencies of the migration clusters.
This completes the proof of our main Theorem, proving that in a suitable sense we expect for any large migration project an expected intrinsic complexity with many small interdependent clusters. ∎
We thus showed that informally we have
Fact.
Any sufficiently large migration project has an intrinsic complexity due to many dependent (comparatively small) migration clusters. The migration of these clusters either takes many successive steps or includes at least one particularly difficult one.
[…]
TABLE I: Numerical evaluation of the expected behavior of migration clusters for migration projects with $n$ components rounded to the nearest integer. We see that one does not expect large mutual dependencies but the expected number of independent clusters is vast and the chain-dependencies are also large when we bound the number of migration clusters with fixed canonical migration time by $f(n)=\mathop{\text{log}}\nolimits n$ or $f(n)=\sqrt{n}/\mathop{\text{log}}\nolimits n$, respectively.
We conclude that even though the model does not represent the exact structure of every migration project, it gives a first sound basis for understanding the structure of migration projects in general. We summarize these results with a numerical evaluation of the relevant estimates in Table I with notation from Appendix -A.
V Practical Implications
Ott et al. define the term cryptographic agility as “the ability to make a transition from one cryptographic algorithm or implementation to another” [18]. Adopting cryptographic agility in that way ensures the flexibility to transition between algorithms or implementations in response to evolving threats. Thus, in order to implement cryptographic agility we need to prepare cryptographic migration. In turn, for preparing migration, we need a step-by-step migration plan. This plan must allow careful re-evaluation of dependencies and system integrity at each stage. This kind of approach is in line with making security solutions sustainable by allowing to adapt to changing threats and technologies. We have shown that the discussed migration approach based on graph models allows step-by-step evaluation and reevaluation throughout the whole process.
Furthermore, we have shown – as stated in Theorem 1 – big migration projects are intrinsically complex. These projects require many successive steps and in-depth strategic planning. Migration graphs can help in understanding this complexity involved and in considering the time needed to migrate, but also specify the number of components and the various dependencies. As we have seen, these dependencies can exist between single components, between interdependent clusters of components, and between external factors such as regulatory requirements, legal issues, or third-party components. The complexity of migrations requires expert knowledge and sufficient dedicated resources, resulting in significant costs. Economic implications and the risks of both, migrating and not migrating to PQC, must be evaluated. It is also crucial for stakeholders, policy makers, and non-technical decision makers to grasp this complexity and to develop appropriate organizational measures to address it. The model we propose can serve as a starting point to identify the highest priority parts of the system, e. g., by identifying the parts which are at the beginning of a migration chain, or which have the highest impact. We expect that from there, an iterative approach works best and we intend to investigate this in future work.
Given this general structure of practical migration, we think that our model can significantly support and enhance existing migration strategies and tools, particularly in the standardization of migration projects through guidelines, rules, and best practices. These models can optimize the process, provide practical examples for guidance, and reduce cost and effort. Initial steps towards such standardization include frameworks like CARAF, a risk management approach for post-quantum migration [12]. CARAF incorporates a threat and a risk assessment stage where our migration graphs can prove valuable. During threat assessment, migration clusters identified using our approach can be used to identify vulnerable subsystems. If one of the systems in the cluster is vulnerable, so is the cluster in its entirety. Likewise, during risk assessment, information about which systems need to be migrated together is helpful to establish cost and effort estimates. Consequently, we believe our approach complements and strengthens existing methods in the field, such as those explored by Näther et al. [14].
Currently, our approach is not supported by tools. We believe that the development and refinement of open-source tools and frameworks, along with ensuring easy access to these resources, can significantly lower the barriers for migration, especially for smaller organizations. These tools can streamline migration processes and offer cost-effective solutions for managing complex projects. We envision tooling to support the creation and maintenance of migration graphs as well as their analysis. Similar tools can be found in related areas such as network analysis and visualization, SBOMs and CBOMs, and graph-based databases. The formalization approach may facilitate the use and adaption of such tools by providing a structured understanding of the underlying complexities. Our future work will also address costs and effort associated with migrations and how these can be captured in the graphs.
In summary, managing the intrinsic complexity of large migration projects involves understanding the dependencies and strategically planning of the different migration steps, while keeping the economic implications in mind. Supporting companies with standardized guidelines and open-source tools is crucial, as is fostering an agile approach to sustainable security solutions. By addressing these key areas, organizations can overcome migration challenges and maintain robust security within their IT-infrastructures.
VI Conclusion
In this work, we used classical mathematical results from graph theory, probability theory, and combinatorial analysis to assess the difficulty of a very practical problem: Cryptographic migration of large IT-infrastructures. We formulated a formal model of cryptographic migratability in terms of directed graphs and showed several expected properties of these graphs. This approach not only elevates the discussion to a scientific level but also establishes a foundation for further research on more efficient migration strategies. In summary, we showed that migration projects always have an intrinsic complexity due to many dependent (comparatively small) migration clusters which (successively) depend on each other.
Still, there is a lot of work to be done, when it comes to practical migration projects: It is not clear how to obtain useful migration graphs in practice, such that they are neither trivial nor too vast. Analyzing how to model real-world IT-infrastructures or relevant subsets will clearly be a challenging task. Also, there is clearly room for improvements: As migration problems typically come with some kind of measurable cost per migration step, a refinement of our model would result in migration graphs where optimal migration strategies are not as easy to obtain as here. One further important open question is also the development of migration graphs and the changes in such graphs over time. Our work sets the stage for future research aimed at refining these models, integrating them with real-world data, and ultimately creating more adaptive, user-friendly tools that support real-world migration. We therefore aim here to scientifically substantiate migration approaches. By emphasizing the relevance of this formal approach, we hope to support the development of more robust migration strategies and (open-source) tooling that can keep pace with the evolving landscape of cryptographic standards. Future work should apply our model to real-world migration projects. The hope is to get a better insight of the complexity such projects not just in general, as discussed here, but for specific practically occurring cases. This might help to better understand real-world cryptographic migration processes, specifically in the context of PQC migration.
Acknowledgments
The authors thank Tobias J. Bauer for his valuable comments during the genesis of this article.
-A Migration Graphs
@@ Acknowledgments
Figure 2: An example of a simple migration graph.
-B Migration Clusters and Migratable Components
@@
In the literature, the migration clusters are classically called the strongly connected components of $G$. For sake of exposition we stick to the term migration cluster in the context of cryptographic migration. Intuitively, the elements of $c(v)$ are precisely those components that have to be migrated at the same time as $v$ due to mutual dependencies. The migration clusters specify how complex single migration steps of a migration project will be. Additionally, the properties of the migration clusters inside a migration graph $G$ will shed light on the complexity of the resulting migration problem which we will describe below. In fact, we have the following important characteristics:
@@
In the case $\mathop{\text{cdeg}}\nolimits(v)=1$ the migration cluster $c(v)$ is a cycle. In practice this case often occurs for $s(v)=2$ when there are two components that communicate using the same protocol. A migration of the protocol can only be done successfully, if both components are migrated at the same time. Cycles of larger sizes are uncommon in practice, cf. Section III-B. The set $m(v)$ of migratable components around $v$ is given by
@@
where we write $\bar{c}(v)=V\setminus c(v)$. The set $m(v)$ specifies exactly which components can be migrated in $G$ when we migrate $v\in V$. If $v$ has dependencies outside its migration cluster, i. e. $v$ cannot be migrated, then $m(v)=\emptyset$. If there are no direct dependencies of $v$, then $m(v)=\left\{v\right\}$. Also, in all other cases $m(v)=c(v)$, i. e. the whole migration cluster. If $m(v)\neq\emptyset$, we call $v$ migratable (in $G$).
Now, in a migration project, there might be several subsequent steps to fully migrate all involved components, i. e. solve the migration problem the migration graph poses. We write $G^{(v)}=(V\setminus m(v),E\setminus(V\to m(v)))$ for the graph after a migration attempt of $v$. If $v$ is not migratable, then $G^{(v)}=G$. Generally, $G^{(v)}$ is exactly the graph restricted to the components outside the migratable components around $v$. Using similar to above the notation $\bar{m}(v)=V\setminus m(v)$, we can thus alternatively write $G^{(v)}=G|_{\bar{m}(v)}=(\bar{m}(v),\bar{m}(v)\to\bar{m}(v))$. We now show
@@
Suppose there is no migratable component, i. e. $m(v)=\emptyset$ for all $v\in V$. By definition, $c(v)\to\bar{c}(v)$ is always non-empty, i. e. there is always a dependency to a component outside the migration cluster of $v$. Call such a dependency $w\in\bar{c}(v)$. Since $w$ is outside the migration cluster of $v$, $w$ cannot depend on elements of $c(v)$ and all dependencies of $w$ reside in the (strict) sub-graph $G|_{\bar{c}(v)}$. Repeating this process, after a finite number of steps, we end-up in a graph which only contains a single migration cluster $c(u)$ and $u\in V$ is migratable in $G$ contradicting our assumption. ∎
-C Migration Strategies
We will now define migration strategies of $G$. Formally, a strategy $S=(v_{1},\ldots,v_{\ell})\in V^{\ell}$ for $\ell\in\mathbb{N}$ is a sequence of components that are migrated one by one (of course together with their respective migration clusters). $\ell$ is the length of the strategy. For $1\leq i\leq\ell$ denote by $G_{i}$ the graph after successively migrating $v_{1},\ldots,v_{i}$, i. e. $G_{i}:=G_{i-1}^{(v_{i})}$, where $G_{0}=G$. The strategy $S$ is reasonable if for all $1\leq i\leq\ell$, we have $m(v_{i})\neq\emptyset$ in the graph $G_{i-1}$. That means that the component $v_{i}$ is in fact migratable in the graph $G_{i-1}$ and, therefore $G_{i}\subsetneq G_{i-1}$, i. e. the migration of $v_{i}$ yields a smaller graph. $S$ is successful, if $G_{\ell}$ is the empty graph. By Lemma 1 there is a reasonable, successful strategy for any migration graph $G$.

### v4

@@ III-A Informal Description of the Model
Also, regulatory requirements of post-quantum solutions and different views on current and future algorithm recommendations may become significantly more complicated, especially in a multinational context. All in all, post-quantum migration is the scenario that motivated our formal migration model. At the same time, it implicates several special properties which make modeling even more complex as we demonstrate in Section V-C.
We now specify the formal model. The concepts concerning directed graphs are partially quite standard, nevertheless we specify for the sake of completeness the full notation. Novel and unprecedented in the literature is the application to migration problems. The reader interested in practical applicability might want to skip the following subsections and the complexity proofs of Section IV, and continue reading in Section V.
III-B Formal Definition of Migration Graphs
@@
Figure 1: An example of a simple migration graph on twelve components $v_{1},\dots,v_{12}$.
III-C Migration Clusters and Migratable Components
@@
In the literature, the migration clusters are classically called the strongly connected components of $G$. For sake of exposition we stick to the term migration cluster in the context of cryptographic migration. Intuitively, the elements of $c(v)$ are precisely those components that have to be migrated at the same time as $v$ due to mutual dependencies. The migration clusters specify how complex single migration steps of a migration project are. Additionally, the properties of the migration clusters inside a migration graph $G$ shed light on the complexity of the resulting migration problem which we will describe below. In fact, we have the following important characteristics:
@@
In the case $\mathop{\text{cdeg}}\nolimits(v)=1$ the migration cluster $c(v)$ is a cycle. In practice this case often occurs for $s(v)=2$ when there are two components that communicate using the same protocol. A migration of the protocol can only be done successfully if both components are migrated at the same time. Cycles of larger sizes are uncommon in practice, cf. Section V-C. The set $m(v)$ of migratable components around $v$ is given by
@@
where we write $\bar{c}(v)=V\setminus c(v)$. The set $m(v)$ specifies exactly which components can be migrated in $G$ when we migrate $v\in V$. If $v$ has dependencies outside its migration cluster, i. e., $v$ cannot be migrated, then $m(v)=\emptyset$. If there are no direct dependencies of $v$, then $m(v)=\left\{v\right\}$. Also, in all other cases $m(v)=c(v)$, i. e., the whole migration cluster. If $m(v)\neq\emptyset$, we call $v$ migratable (in $G$).
Now, in a migration project, there might be several subsequent steps to fully migrate all involved components, i. e., solve the migration problem the migration graph poses. We write $G^{(v)}=(V\setminus m(v),E\setminus(V\to m(v)))$ for the graph after a migration attempt of $v$. If $v$ is not migratable, then $G^{(v)}=G$. Generally, $G^{(v)}$ is exactly the graph restricted to the components outside the migratable components around $v$. Using similar to above the notation $\bar{m}(v)=V\setminus m(v)$, we can thus alternatively write $G^{(v)}=G|_{\bar{m}(v)}=(\bar{m}(v),\bar{m}(v)\to\bar{m}(v))$. We now show
@@
Suppose there is no migratable component, i. e., $m(v)=\emptyset$ for all $v\in V$. By definition, $c(v)\to\bar{c}(v)$ is always non-empty, i. e., there is always a dependency to a component outside the migration cluster of $v$. Call such a dependency $w\in\bar{c}(v)$. Since $w$ is outside the migration cluster of $v$, $w$ cannot depend on elements of $c(v)$ and all dependencies of $w$ reside in the (strict) sub-graph $G|_{\bar{c}(v)}$. Repeating this process, after a finite number of steps, we end-up in a graph which only contains a single migration cluster $c(u)$ and $u\in V$ is migratable in $G$ contradicting our assumption. ∎
III-D Migration Strategies
We will now define migration strategies of $G$. Formally, a strategy $S=(v_{1},\ldots,v_{\ell})\in V^{\ell}$ for $\ell\in\mathbb{N}$ is a sequence of components that are migrated one by one (of course together with their respective migration clusters). $\ell$ is the length of the strategy. For $1\leq i\leq\ell$ denote by $G_{i}$ the graph after successively migrating $v_{1},\ldots,v_{i}$, i. e., $G_{i}:=G_{i-1}^{(v_{i})}$, where $G_{0}=G$. The strategy $S$ is reasonable if for all $1\leq i\leq\ell$, we have $m(v_{i})\neq\emptyset$ in the graph $G_{i-1}$. That means that the component $v_{i}$ is in fact migratable in the graph $G_{i-1}$ and, therefore $G_{i}\subsetneq G_{i-1}$, i. e., the migration of $v_{i}$ yields a smaller graph. $S$ is successful, if $G_{\ell}$ is the empty graph. By Lemma 1 there is a reasonable, successful strategy for any migration graph $G$.
$c_{7}$$c_{8}$$c_{9}$$c_{6}$$c_{5}$$c_{4}$$c_{2}$$c_{3}$$c_{1}$
Figure 2: The acyclic condensation graph $G/c$ on nine clusters $c_{1},\dots,c_{9}$ of the migration graph from Fig. 1.